OpenStudy (anonymous):

factor to find zeros of y=x^2+4x-21

OpenStudy (anonymous):

for its zeros, value of y=0 so x^2 + 4x - 21 = 0 x^2 + 7x - 3x - 21 = 0 x(x + 7) -3(x + 7) = 0 (x - 3)(X + 7) = 0 if x - 3 = 0, then x = 3 if x + 7 = 0, then x = -7 so the zeros of y are 3 and -7